# Restricted $L_{\infty}$-algebras and a derived Milnor-Moore theorem

### v1

@@ 7.1. Rational stable $\infty$-categories
Since ${\mathrm{Mod}}_{{\mathrm{H}}({\mathbb{Q}})}(\mathrm{Sp})$ is stable, by [12, Theorem 4.6.] the canonical map
@@
identifies with induction along the canonical map of spectra ${\mathrm{H}}({\mathbb{Q}})\simeq{\mathbb{1}}\otimes{\mathrm{H}}({\mathbb{Q}})\to{\mathrm{H}}({\mathbb{Q}})\otimes{\mathrm{H}}({\mathbb{Q}}),$ which is an equivalence. By [23, Proposition 4.8.2.10.] this guarantees that the forgetful functor ${\mathrm{Mod}}_{{\mathrm{Mod}}_{{\mathrm{H}}({\mathbb{Q}})}(\mathrm{Sp})}(\mathrm{Pr}^{\mathrm{L}})\to\mathrm{Pr}^{\mathrm{L}}$ is fully faithful. We characterize the essential image.
Every left module in $\mathrm{Pr}^{\mathrm{L}}$ over ${\mathrm{Mod}}_{{\mathrm{H}}({\mathbb{Q}})}(\mathrm{Sp})$ has an underlying left module over $\mathrm{Sp}$ forgetting along the symmetric monoidal left adjoint functor $\mathrm{Sp}\to{\mathrm{Mod}}_{{\mathrm{H}}({\mathbb{Q}})}(\mathrm{Sp})$. By [12, Theorem 4.6.] every left module over $\mathrm{Sp}$ in $\mathrm{Pr}^{\mathrm{L}}$ is a stable $\infty$-category, and for every stable presentable $\infty$-category ${\mathcal{B}}$ the canonical map ${\mathcal{B}}\simeq{\mathcal{S}}\otimes{\mathcal{B}}\to\mathrm{Sp}\otimes{\mathcal{B}}$ is an equivalence. Consequently, it is enough to see that for any stable presentable $\infty$-category ${\mathcal{C}}$ the following holds: the hom sets of the homotopy category of ${\mathcal{C}}$ are ${\mathbb{Q}}$-vector spaces if and only if the canonical map ${\mathcal{C}}\simeq{\mathcal{S}}\otimes{\mathcal{C}}\to{\mathrm{Mod}}_{{\mathrm{H}}({\mathbb{Q}})}(\mathrm{Sp})\otimes{\mathcal{C}}$ is an equivalence. The latter map identifies with the functor
@@
where the last equivalence is by [23, Theorem 4.8.4.6.]. The latter functor identifies with the free ${\mathrm{H}}({\mathbb{Q}})$-module functor. Since the canonical map of spectra ${\mathrm{H}}({\mathbb{Q}})\simeq{\mathbb{1}}\otimes{\mathrm{H}}({\mathbb{Q}})\to{\mathrm{H}}({\mathbb{Q}})\otimes{\mathrm{H}}({\mathbb{Q}})$ is an equivalence, by [23, Proposition 4.8.2.10.] the forgetful functor ${\mathrm{Mod}}_{{\mathrm{H}}({\mathbb{Q}})}({\mathcal{C}})\to{\mathcal{C}}$ is fully faithful and the essential image consists of those ${\mathrm{X}}\in{\mathcal{C}}$ such that one of the following equivalent conditions hold:
@@ 7.2. ${\mathcal{O}}$-monoidal closures
For the next notation we use that for every small ${\mathcal{O}}$-monoidal $\infty$-category ${\mathcal{C}}^{\otimes}\to{\mathcal{O}}^{\otimes}$ there is an ${\mathcal{O}}$-monoidal Yoneda-embedding ${\mathcal{C}}^{\otimes}\to{\mathcal{P}}({\mathcal{C}})^{\otimes}$ [23, Proposition 4.8.1.10. ].
Notation 7.4.
Let ${\mathcal{K}},{\mathcal{K}}^{\prime}\subset\mathsf{Cat}_{\infty}$ be full subcategories and ${\mathcal{C}}^{\otimes}\to{\mathcal{O}}^{\otimes}$ an ${\mathcal{O}}$-monoidal $\infty$-category. Let
@@
be the full suboperad spanned by the functors ${\mathcal{C}}_{\mathrm{X}}^{\mathrm{op}}\to\widehat{{\mathcal{S}}}$ preserving ${\mathcal{K}}$-indexed limits for some ${\mathrm{X}}\in{\mathcal{O}}$. Let
@@
be the full suboperad such that for every ${\mathrm{X}}\in{\mathcal{O}}$ the full subcategory ${\mathcal{P}}^{{\mathcal{K}}^{\prime}}_{\mathcal{K}}({\mathcal{C}})_{\mathrm{X}}\subset\widehat{{\mathcal{P}}}_{\mathcal{K}}({\mathcal{C}})_{\mathrm{X}}$ is generated by ${\mathcal{C}}_{\mathrm{X}}$ under ${\mathcal{K}}^{\prime}$-indexed colimits.
@@
If ${\mathcal{K}}=\mathsf{Cat}_{\infty}$, we have ${\mathcal{P}}^{{\mathcal{K}}^{\prime}}_{\mathcal{K}}({\mathcal{C}})^{\otimes}={\mathcal{C}}^{\otimes}.$
@@
Let ${\mathcal{K}},{\mathcal{K}}^{\prime}\subset\mathsf{Cat}_{\infty}$ be full subcategories and ${\mathcal{C}}^{\otimes}\to{\mathcal{O}}^{\otimes}$ an ${\mathcal{O}}$-monoidal $\infty$-category.
@@
    Let ${\mathrm{X}}\in{\mathcal{O}}$. If the fiber ${\mathcal{C}}_{\mathrm{X}}$ has ${\mathcal{K}}$-indexed colimits, $\widehat{{\mathcal{P}}}_{\mathcal{K}}({\mathcal{C}})_{\mathrm{X}}^{\otimes}\subset\widehat{{\mathcal{P}}}({\mathcal{C}})^{\otimes}_{\mathrm{X}}$ is a localization. In particular, $\widehat{{\mathcal{P}}}_{\mathcal{K}}({\mathcal{C}})^{\otimes}_{\mathrm{X}}$ has large colimits and so ${\mathcal{P}}^{{\mathcal{K}}^{\prime}}_{\mathcal{K}}({\mathcal{C}})^{\otimes}_{\mathrm{X}}$ has ${\mathcal{K}}^{\prime}$-indexed colimits.
@@
    If the ${\mathcal{O}}$-monoidal $\infty$-category ${\mathcal{C}}^{\otimes}\to{\mathcal{O}}^{\otimes}$ is compatible with ${\mathcal{K}}$-indexed colimits, $\widehat{{\mathcal{P}}}_{\mathcal{K}}({\mathcal{C}})^{\otimes}\subset\widehat{{\mathcal{P}}}({\mathcal{C}})^{\otimes}$ is a localization relative to ${\mathcal{O}}^{\otimes}$. In particular, $\widehat{{\mathcal{P}}}_{\mathcal{K}}({\mathcal{C}})^{\otimes}\to{\mathcal{O}}^{\otimes}$ is an ${\mathcal{O}}$-monoidal $\infty$-category compatible with large colimits and so ${\mathcal{P}}^{{\mathcal{K}}^{\prime}}_{\mathcal{K}}({\mathcal{C}})^{\otimes}\to{\mathcal{O}}^{\otimes}$ is an ${\mathcal{O}}$-monoidal $\infty$-category compatible with ${\mathcal{K}}^{\prime}$-indexed colimits.
@@
Let ${\mathcal{K}},{\mathcal{K}}^{\prime}\subset\mathsf{Cat}_{\infty}$ be full subcategories and ${\mathcal{C}}^{\otimes}\to{\mathcal{O}}^{\otimes}$ an ${\mathcal{O}}$-monoidal $\infty$-category. Let ${\mathrm{X}}\in{\mathcal{O}}$. If the fiber ${\mathcal{C}}_{\mathrm{X}}$ has ${\mathcal{K}}^{\prime}$-indexed colimits, ${\mathcal{C}}_{\mathrm{X}}^{\otimes}\subset{\mathcal{P}}^{{\mathcal{K}}^{\prime}}_{\mathcal{K}}({\mathcal{C}})_{\mathrm{X}}^{\otimes}$ is a localization.
@@
First assume that ${\mathcal{C}}_{\mathrm{X}}$ has small colimits. Then by Lemma 7.7 (1) the embeddings $\widehat{{\mathcal{P}}}_{\mathsf{Cat}_{\infty}}({\mathcal{C}})_{\mathrm{X}}\subset\widehat{{\mathcal{P}}}({\mathcal{C}})_{\mathrm{X}},\widehat{{\mathcal{P}}}_{\mathcal{K}}({\mathcal{C}})_{\mathrm{X}}\subset\widehat{{\mathcal{P}}}({\mathcal{C}})_{\mathrm{X}}$ admit a left adjoint. Hence the embedding $\widehat{{\mathcal{P}}}_{\mathsf{Cat}_{\infty}}({\mathcal{C}})_{\mathrm{X}}\subset\widehat{{\mathcal{P}}}_{\mathcal{K}}({\mathcal{C}})_{\mathrm{X}}$ admits a left adjoint. Thus the embedding ${\mathcal{C}}_{\mathrm{X}}={\mathcal{P}}^{{\mathcal{K}}^{\prime}}_{\mathsf{Cat}_{\infty}}({\mathcal{C}})_{\mathrm{X}}\subset{\mathcal{P}}^{{\mathcal{K}}^{\prime}}_{\mathcal{K}}({\mathcal{C}})_{\mathrm{X}}$ admits a left adjoint.
In the general case we set ${\mathcal{D}}^{\otimes}:={\mathcal{P}}^{\mathsf{Cat}_{\infty}}_{{\mathcal{K}}^{\prime}}({\mathcal{C}})^{\otimes}$. By Lemma 7.7 (1) the $\infty$-category ${\mathcal{D}}_{\mathrm{X}}$ admits small colimits. Moreover the embedding ${\mathcal{C}}_{\mathrm{X}}\subset{\mathcal{D}}_{\mathrm{X}}$ preserves ${\mathcal{K}}^{\prime}$-indexed colimits. The first part of the proof shows that the embedding ${\mathcal{D}}_{\mathrm{X}}\subset{\mathcal{P}}^{{\mathcal{K}}^{\prime}}_{\mathcal{K}}({\mathcal{D}})_{\mathrm{X}}$ admits a left adjoint. The left adjoint restricts to a functor ${\mathcal{P}}^{{\mathcal{K}}^{\prime}}_{\mathcal{K}}({\mathcal{C}})_{\mathrm{X}}\to{\mathcal{C}}_{\mathrm{X}}$ since ${\mathcal{C}}_{\mathrm{X}}$ is closed in ${\mathcal{D}}_{\mathrm{X}}$ under ${\mathcal{K}}^{\prime}$-indexed colimits. Hence the embedding ${\mathcal{C}}_{\mathrm{X}}\subset{\mathcal{P}}^{{\mathcal{K}}^{\prime}}_{\mathcal{K}}({\mathcal{C}})_{\mathrm{X}}$ admits a left adjoint.
@@
Let ${\mathcal{K}},{\mathcal{K}}^{\prime}\subset\mathsf{Cat}_{\infty}$ be full subcategories and ${\mathcal{C}}^{\otimes}\to{\mathcal{O}}^{\otimes}$ an ${\mathcal{O}}$-monoidal $\infty$-category compatible with ${\mathcal{K}}$-indexed colimits that admits fiberwise ${\mathcal{K}}^{\prime}$-indexed colimits.
@@
  • ${\mathcal{D}}^{\otimes}\to{\mathcal{O}}^{\otimes}$ is compatible with ${\mathcal{K}}^{\prime}$-indexed colimits,
  • for every ${\mathrm{X}}\in{\mathcal{O}}$ the fiber ${\mathcal{D}}_{\mathrm{X}}$ is generated in ${\mathcal{E}}_{\mathrm{X}}$ by ${\mathcal{C}}_{\mathrm{X}}$ under ${\mathcal{K}}^{\prime}$-indexed colimits,
@@
Let ${\mathcal{C}}^{\otimes}\to{\mathcal{O}}^{\otimes}$ be a pointed, preadditive, additive, stable ${\mathcal{O}}$-monoidal $\infty$-category, respectively, that admits fiberwise ${\mathcal{K}}^{\prime}$-indexed colimits, where ${\mathcal{K}}^{\prime}\subset\mathsf{Cat}_{\infty}$ is a full subcategory containing {}$\{\emptyset\},{\mathrm{{\mathcal{F}}in}},\mathsf{Cat}_{\infty}^{\mathrm{fin}}$, respectively, depending on the case.
@@
  • ${\mathcal{D}}^{\otimes}\to{\mathcal{O}}^{\otimes}$ is compatible with ${\mathcal{K}}^{\prime}$-indexed colimits,
  • for every ${\mathrm{X}}\in{\mathcal{O}}$ the fiber ${\mathcal{D}}_{\mathrm{X}}$ is generated in ${\mathcal{E}}_{\mathrm{X}}$ by ${\mathcal{C}}_{\mathrm{X}}$ under ${\mathcal{K}}^{\prime}$-indexed colimits,
@@
Let ${\mathrm{X}}\in{\mathcal{O}}.$ For ${\mathcal{K}}=\{\emptyset\}$ the $\infty$-category $\widehat{{\mathcal{P}}}_{\mathcal{K}}({\mathcal{C}})_{\mathrm{X}}=\mathrm{Fun}_{*}({\mathcal{C}}_{\mathrm{X}}^{\mathrm{op}},\widehat{{\mathcal{S}}})$ of final object preserving functors admits a zero object if ${\mathcal{C}}$ admits a final object. The zero object belongs to the full subcategory ${\mathcal{P}}^{{\mathcal{K}}^{\prime}}_{\mathcal{K}}({\mathcal{C}})$. For ${\mathcal{K}}={\mathrm{{\mathcal{F}}in}}$ the $\infty$-category $\widehat{{\mathcal{P}}}_{\mathcal{K}}({\mathcal{C}})_{\mathrm{X}}=\mathrm{Fun}^{\Pi}({\mathcal{C}}_{\mathrm{X}}^{\mathrm{op}},\widehat{{\mathcal{S}}})$ of finite products preserving functors is (pre)additive if ${\mathcal{C}}$ is (pre)additive [12, Corollary 2.10.]. So the full subcategory ${\mathcal{P}}^{{\mathcal{K}}^{\prime}}_{\mathcal{K}}({\mathcal{C}})$ closed under finite coproducts is (pre)additive. For ${\mathcal{K}}=\mathsf{Cat}_{\infty}^{\mathrm{fin}}$ the $\infty$-category $\widehat{{\mathcal{P}}}_{\mathcal{K}}({\mathcal{C}})=\mathrm{Fun}^{\mathrm{lex}}({\mathcal{C}}^{\mathrm{op}},\widehat{{\mathcal{S}}})$ of finite limits preserving functors is stable if ${\mathcal{C}}$ is stable. So the full subcategory ${\mathcal{P}}^{{\mathcal{K}}^{\prime}}_{\mathcal{K}}({\mathcal{C}})$ closed under finite colimits and arbitrary shifts is stable.
Lemma 7.11.
Let $\phi:{\mathcal{C}}\to{\mathcal{D}}$ be a locally cocartesian fibration and ${\mathcal{B}}\subset{\mathcal{C}}$ a full subcategory such that for every ${\mathrm{Z}}\in{\mathcal{D}}$ the induced embedding on fibers ${\mathcal{B}}_{\mathrm{Z}}\subset{\mathcal{C}}_{\mathrm{Z}}$ admits a left adjoint. The restriction $\phi^{\prime}:{\mathcal{B}}\subset{\mathcal{C}}\to{\mathcal{D}}$ is a locally cocartesian fibration, where a morphism ${\mathrm{X}}\to{\mathrm{Z}}$ in ${\mathcal{B}}$ is locally $\phi^{\prime}$-cocartesian if and only if it factors as a locally $\phi$-cocartesian morphism ${\mathrm{X}}\to{\mathrm{Y}}$ in ${\mathcal{C}}$ followed by a local equivalence ${\mathrm{Y}}\to{\mathrm{Z}}$ of the fiber over $\phi({\mathrm{Z}}).$
Proof.
Since $\phi$ is a locally cocartesian fibration, it will be enough to see that for any locally $\phi$-cocartesian morphism $\alpha:{\mathrm{X}}\to{\mathrm{Y}}$ in ${\mathcal{C}}$ and local equivalence $\beta:{\mathrm{Y}}\to{\mathrm{Z}}$ of the fiber over $\phi({\mathrm{Z}})$ the composition $\beta\circ\alpha:{\mathrm{X}}\to{\mathrm{Y}}\to{\mathrm{Z}}$ is locally $\phi^{\prime}$-cocartesian. We like to see that for any ${\mathrm{T}}\in{\mathcal{B}}_{\phi({\mathrm{Z}})}$ the canonical map
${\mathcal{B}}_{\phi({\mathrm{Z}})}({\mathrm{Z}},{\mathrm{T}})\to\{\phi(\alpha)\}\times_{{\mathcal{D}}(\phi({\mathrm{X}}),\phi({\mathrm{Z}}))}{\mathcal{B}}({\mathrm{X}},{\mathrm{T}})$
is an equivalence. The latter map factors as
${\mathcal{B}}_{\phi({\mathrm{Z}})}({\mathrm{Z}},{\mathrm{T}})\simeq{\mathcal{C}}_{\phi({\mathrm{Z}})}({\mathrm{Y}},{\mathrm{T}})\simeq\{\phi(\alpha)\}\times_{{\mathcal{D}}(\phi({\mathrm{X}}),\phi({\mathrm{Z}}))}{\mathcal{C}}({\mathrm{X}},{\mathrm{T}}).$
Lemma 7.12.
Let ${\mathcal{C}}\to\mathsf{Set}$ be an algebraic functor and ${\mathcal{D}}\to{\mathcal{C}}$ a monadic functor that preserves filtered colimits. Assume that ${\mathcal{C}}$ is generated by the essential image of the free functor ${\mathrm{F}}:\mathsf{Set}\to{\mathcal{C}}$ under small filtered colimits. The functor ${\mathcal{D}}\to{\mathcal{C}}\to\mathsf{Set}$ is algebraic. In particular, the functor ${\mathcal{D}}\to{\mathcal{C}}$ preserves small sifted colimits.
Proof.
Let ${\mathcal{C}}\to\mathsf{Set},{\mathcal{D}}\to{\mathcal{C}}\to\mathsf{Set}$ be the right adjoint functors with associated theories $\mathbb{T},\mathbb{T}^{\prime}.$ The adjunction ${\mathcal{F}}:{\mathcal{C}}\rightleftarrows{\mathcal{D}}:{\mathcal{G}}$ gives rise to a map of theories $\phi:\mathbb{T}\to\mathbb{T}^{\prime}$ that is the restriction of the functor ${\mathcal{F}}^{\mathrm{op}}:{\mathcal{C}}^{\mathrm{op}}\to{\mathcal{D}}^{\mathrm{op}}$. The functor $\phi$ yields an adjunction $\phi_{\ast}:\mathrm{Alg}_{\mathbb{T}}(\mathsf{Set})\rightleftarrows\mathrm{Alg}_{\mathbb{T}^{\prime}}(\mathsf{Set}):\phi^{\ast},$ where the right adjoint is monadic. There is a commutative square
@@
such that the bottom functor is an equivalence since ${\mathcal{C}}\to\mathsf{Set}$ is algebraic and both vertical functors are monadic. Consequently, it is enough to see that the natural transformation
$\alpha:\phi_{\ast}\circ\theta\to\Phi\circ{\mathcal{F}}$
corresponding to the natural transformation $\theta\to\theta\circ{\mathcal{G}}\circ{\mathcal{F}}\cong\phi^{\ast}\circ\Phi\circ{\mathcal{F}}$ is an isomorphism. The functor ${\mathcal{D}}\to{\mathcal{C}}\to\mathsf{Set}$ preserves small filtered colimits so that its left adjoint preserves compact objects. Thus $\mathbb{T}^{\prime\mathrm{op}}\subset{\mathcal{D}}$ consists of compact objects of ${\mathcal{D}}.$ Therefore for every ${\mathrm{X}}\in\mathbb{T}^{\prime}$ the composition ${\mathcal{D}}({\mathrm{X}},-):{\mathcal{D}}\xrightarrow{\Phi}\mathrm{Alg}_{\mathbb{T}^{\prime}}(\mathsf{Set})\xrightarrow{{\mathrm{ev}}_{\mathrm{X}}}\mathsf{Set}$ preserves small filtered colimits so that the functor $\Phi:{\mathcal{D}}\to\mathrm{Alg}_{\mathbb{T}^{\prime}}(\mathsf{Set})$ preserves small filtered colimits. Consequently, by assumption on ${\mathcal{C}}$ it is enough to see that $\alpha({\mathrm{Y}}):\phi_{\ast}(\theta({\mathrm{Y}}))\to\Phi({\mathcal{F}}({\mathrm{Y}}))$ is an isomorphism for every ${\mathrm{Y}}\in{\mathcal{C}}$ that is the image of a finite set under the free functor ${\mathrm{F}}:\mathsf{Set}\to{\mathcal{C}}$.
Let ${\mathrm{H}}\in\mathrm{Alg}_{\mathbb{T}^{\prime}}(\mathsf{Set})$ and ${\mathrm{X}}\in{\mathrm{{\mathcal{F}}in}}.$ Then the induced map
$\mathrm{Alg}_{\mathbb{T}^{\prime}}(\mathsf{Set})(\Phi({\mathcal{F}}({\mathrm{F}}({\mathrm{X}}))),{\mathrm{H}})\to\mathrm{Alg}_{\mathbb{T}^{\prime}}(\mathsf{Set})(\phi_{\ast}(\theta({\mathrm{F}}({\mathrm{X}}))),{\mathrm{H}})$
coincides with the map
$\mathrm{Alg}_{\mathbb{T}^{\prime}}(\mathsf{Set})(\Phi({\mathcal{F}}({\mathrm{F}}({\mathrm{X}}))),{\mathrm{H}})\cong\mathrm{Alg}_{\mathbb{T}^{\prime}}(\mathsf{Set})({\mathrm{T}}^{\prime}({\mathcal{F}}({\mathrm{F}}({\mathrm{X}})),-),{\mathrm{H}})\cong$
$\cong{\mathrm{H}}({\mathcal{F}}({\mathrm{F}}({\mathrm{X}})))\cong{\mathrm{H}}(\phi({\mathrm{F}}({\mathrm{X}})))\cong\mathrm{Alg}_{\mathbb{T}}(\mathsf{Set})({\mathrm{T}}({\mathrm{F}}({\mathrm{X}}),-),\phi^{\ast}({\mathrm{H}}))$
$\cong\mathrm{Alg}_{\mathbb{T}}(\mathsf{Set})(\theta({\mathrm{F}}({\mathrm{X}})),\phi^{\ast}({\mathrm{H}}))\cong\mathrm{Alg}_{\mathbb{T}^{\prime}}(\mathsf{Set})(\phi_{\ast}(\theta({\mathrm{F}}({\mathrm{X}}))),{\mathrm{H}}).$

### v2

@@ 5.1. Rational stable $\infty$-categories
Since ${\mathrm{Mod}}_{{\mathrm{H}}({\mathbb{Q}})}({\mathrm{Sp}})$ is stable, by [19, Theorem 4.6.] the canonical map
@@
identifies with induction along the canonical map of spectra ${\mathrm{H}}({\mathbb{Q}})\simeq{\mathbb{1}}\otimes{\mathrm{H}}({\mathbb{Q}})\to{\mathrm{H}}({\mathbb{Q}})\otimes{\mathrm{H}}({\mathbb{Q}}),$ which is an equivalence. By [39, Proposition 4.8.2.10.] this guarantees that the forgetful functor ${\mathrm{Mod}}_{{\mathrm{Mod}}_{{\mathrm{H}}({\mathbb{Q}})}({\mathrm{Sp}})}(\mathrm{Pr}^{\mathrm{L}})\to\mathrm{Pr}^{\mathrm{L}}$ is fully faithful. We characterize the essential image.
Every left module in $\mathrm{Pr}^{\mathrm{L}}$ over ${\mathrm{Mod}}_{{\mathrm{H}}({\mathbb{Q}})}({\mathrm{Sp}})$ has an underlying left module over $\mathrm{Sp}$} forgetting along the symmetric monoidal left adjoint functor ${\mathrm{Sp}}\to{\mathrm{Mod}}_{{\mathrm{H}}({\mathbb{Q}})}({\mathrm{Sp}})$. By [26, Corollary 9.4.] every left module over $\mathrm{Sp}$} in $\mathrm{Pr}^{\mathrm{L}}$ is a stable $\infty$-category, and for every stable presentable $\infty$-category $\mathcal{B}$} the canonical map ${\mathcal{B}}\simeq{\mathcal{S}}\otimes{\mathcal{B}}\to{\mathrm{Sp}}\otimes{\mathcal{B}}$ is an equivalence. Consequently, it is enough to see that for any stable presentable $\infty$-category $\mathcal{C}$} the following holds: the hom sets of the homotopy category of $\mathcal{C}$} are $\mathbb{Q}$}-vector spaces if and only if the canonical map ${\mathcal{C}}\simeq{\mathcal{S}}\otimes{\mathcal{C}}\to{\mathrm{Mod}}_{{\mathrm{H}}({\mathbb{Q}})}({\mathrm{Sp}})\otimes{\mathcal{C}}$ is an equivalence. The latter map identifies with the functor
@@
where the last equivalence is by [39, Theorem 4.8.4.6.]. The latter functor identifies with the free ${\mathrm{H}}({\mathbb{Q}})$-module functor. Since the canonical map of spectra ${\mathrm{H}}({\mathbb{Q}})\simeq{\mathbb{1}}\otimes{\mathrm{H}}({\mathbb{Q}})\to{\mathrm{H}}({\mathbb{Q}})\otimes{\mathrm{H}}({\mathbb{Q}})$ is an equivalence, by [39, Proposition 4.8.2.10.] the forgetful functor ${\mathrm{Mod}}_{{\mathrm{H}}({\mathbb{Q}})}({\mathcal{C}})\to{\mathcal{C}}$ is fully faithful and the essential image consists of those ${\mathrm{X}}\in{\mathcal{C}}$ such that one of the following equivalent conditions hold:
@@ 5.2. $\mathcal{O}$}-monoidal closures
For the next notation we use that for every small $\mathcal{O}$}-monoidal $\infty$-category ${\mathcal{C}}^{\otimes}\to{\mathcal{O}}^{\otimes}$ there is an $\mathcal{O}$}-monoidal Yoneda-embedding ${\mathcal{C}}^{\otimes}\to{\mathcal{P}}({\mathcal{C}})^{\otimes}$ [39, Proposition 4.8.1.10. ].
Notation 5.4.
Let ${\mathcal{K}},{\mathcal{K}}^{\prime}\subset\mathrm{Cat}_{\infty}$ be full subcategories and ${\mathcal{C}}^{\otimes}\to{\mathcal{O}}^{\otimes}$ an $\mathcal{O}$}-monoidal $\infty$-category. Let
@@
be the full suboperad spanned by the functors $^{\mathcal{C}}_{\mathrm{X}}^{\mathrm{op}}\to\widehat{{\mathcal{S}}}$ preserving $\mathcal{K}$}-indexed limits for some ${\mathrm{X}}\in{\mathcal{O}}$. Let
@@
[…]
@@
If ${\mathcal{K}}=\mathrm{Cat}_{\infty}$, we have ${\mathcal{P}}^{{\mathcal{K}}^{\prime}}_{\mathcal{K}}({\mathcal{C}})^{\otimes}={\mathcal{C}}^{\otimes}.$
@@
Let ${\mathcal{K}},{\mathcal{K}}^{\prime}\subset\mathrm{Cat}_{\infty}$ be full subcategories and ${\mathcal{C}}^{\otimes}\to{\mathcal{O}}^{\otimes}$ an $\mathcal{O}$}-monoidal $\infty$-category.
@@
[…]
@@
[…]
@@
[…]
@@
First assume that ${\mathcal{C}}_{\mathrm{X}}$ has small colimits. Then by Lemma 5.7 (1) the embeddings $\widehat{{\mathcal{P}}}_{\mathrm{Cat}_{\infty}}({\mathcal{C}})_{\mathrm{X}}\subset\widehat{{\mathcal{P}}}({\mathcal{C}})_{\mathrm{X}},\widehat{{\mathcal{P}}}_{\mathcal{K}}({\mathcal{C}})_{\mathrm{X}}\subset\widehat{{\mathcal{P}}}({\mathcal{C}})_{\mathrm{X}}$ admit a left adjoint. Hence the embedding $\widehat{{\mathcal{P}}}_{\mathrm{Cat}_{\infty}}({\mathcal{C}})_{\mathrm{X}}\subset\widehat{{\mathcal{P}}}_{\mathcal{K}}({\mathcal{C}})_{\mathrm{X}}$ admits a left adjoint. Thus the embedding ${\mathcal{C}}_{\mathrm{X}}={\mathcal{P}}^{{\mathcal{K}}^{\prime}}_{\mathrm{Cat}_{\infty}}({\mathcal{C}})_{\mathrm{X}}\subset{\mathcal{P}}^{{\mathcal{K}}^{\prime}}_{\mathcal{K}}({\mathcal{C}})_{\mathrm{X}}$ admits a left adjoint.
[…]
@@
[…]
@@
[…]
@@
Notation 5.10.
Let $\mathrm{Cat}_{\infty}^{\mathrm{fin}}\subset\mathrm{Cat}_{\infty}$ be the full subcategory generated by the final $\infty$-category under finite colimits.
@@
[…]
@@
[…]